Cawley Company makes three models of tasers. Information on the three products is given below Sales Variable expenses Contribution margin Fixed expenses Net income Tingler Shocker Stunner $300,000 S500,000 $200,000 135,000 65,000 93,000 $18,000 $85,000 (28,000) 165,000 190,000 135,000 117,000 225,000 310,000 Fixed expenses consist of $290,000 of common costs allocated to the three products based on relative sales, and additional fixed expenses of $30,000 (Tingler), $80,000 (Shocker), and $35,000 (Stunner). The common costs will be incurred regardless of how many models are produced. The other fixed expenses would be eliminated if a model is phased out. James Watt, an executive with the company, feels the Stunner line should be discontinued to increase the companys net income. Compute current net income for Cawley Company Net income s Compute net income by product line and in total for Cawley Company if the company discontinues the Stunner product line. (Hint: Allocate the $290,000 common costs to the two remaining product lines based on thein relative sales.) (Round intermediate calculations to 4 decimal places, e.g. 15.2575 and final answers to the nearest whole dollar, e.g. 5,275.) Tingler Net Income Shocker Net Income Total Net Income
Answer #1
a
Net income 75000
b
Tingler Net Income -3750
Shocker Net Income 48750
Total Net Income 45000
Workings:
a
Tingler net income 18000
Shocker net income 85000
Stunner net income -28000
Net income 75000
b
Tingler Shocker Total
Sales 300000 500000 800000
Variable expenses 165000 190000 355000
Contribution margin 135000 310000 445000
Fixed expenses 138750 261250 400000
Net income -3750 48750 45000
Fixed expenses :
Tingler= 30000+(290000/800000*300000) = $138750
Shocker= 80000+(290000/800000*500000) = $261250